OpenStudy (anonymous):

Simplify the expression. 2^2 • 2^3 A. 4^5 B. 4^6 C. 2^6 D. 2^5

OpenStudy (anonymous):

a

OpenStudy (anonymous):

d

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Trust me its A

OpenStudy (anonymous):

just plus the exponent 2+3 = 5

OpenStudy (anonymous):

what is the real answer?

OpenStudy (anonymous):

no, it is D 2^5=32

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

NO! YOUR WRONG ITS A

OpenStudy (anonymous):

????

OpenStudy (anonymous):

JINX U WILL GET IT WRONG

OpenStudy (anonymous):

why dont you simplied 2^2=4 and 2^3=8, so 4*8=32

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

ALRIGHT I WILL TAKE YOU THROUGH IT STEP BY STEP OK

OpenStudy (anonymous):

im confused....redneck, show me how you got 4^5

OpenStudy (anonymous):

The redneck is messing with you, it's D.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

2^2 * 2^3 OK SO YOU TAKE THE EXPONENTS AND ADD NOT ONLY ARE YOU ADDING YOU ARE MULTIPLY THE ROOT WHICH IS 2 SO 2*2 = 4 AND THE EXPONENTS WERE ADDED SO THAT MAKE THE ANSWER 4^5

OpenStudy (anonymous):

When you have a^(c) and you multiply that by a^(d) the general rule is that it is equal to a^(c+d)

OpenStudy (anonymous):

YOUR FORGETTING THE ROOT

OpenStudy (anonymous):

no, you dont need to multiply the root

OpenStudy (anonymous):

I AM 17 AND DO CALCULAS IM PRETTY SURE I KNOW WHAT THE HELL TO DO OK

OpenStudy (anonymous):

Okay man, if that's the case, I'm 20, in university and I have been in top 20 in maths in my country.

OpenStudy (anonymous):

what to do....?

OpenStudy (anonymous):

if you don't trust me, click here http://www.wolframalpha.com/input/?i=2%5E2*2%5E3

OpenStudy (anonymous):

why dont you check that in your caculator, you will get your answer 32

OpenStudy (anonymous):

HOLD ON ONE MINUTE OK YOU ALL HAVE A FREE SHOT ON ME DO TO MY STUPIDITY I WAS WRONG CHENG THANKS FOR TELLING ME I AM USED TO BIGGER PROBLEMS WITH FOIL

OpenStudy (anonymous):

D IS THE RIGHT ANSWER IM SORRY

OpenStudy (anonymous):

FOIL HAS A DIFFERENT RULE THATS WHY I HAD THOUGHT IT WAS A

OpenStudy (anonymous):

ok thanks redneck4302 @ashleycheng123 and @dreamer02. ps seriously @deavol you're reporting rednck just for not being sure how to solve a problem? I will report YOU now

myininaya (myininaya):

Mistakes happen. Sometimes are logic gets confused. It is alright. This happens. We are studying together to get better. We are all learning.

myininaya (myininaya):

Like just now. I said are instead of our.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok I'll close the question now ok

OpenStudy (anonymous):

@myininaya, I saw his posts in other topics as well, they were completely random answers.

OpenStudy (anonymous):

how long do i have to wait for aravinds reply?

OpenStudy (aravindg):

Yeah but I wanted to remind him of 2 things: 1.It is better not to use CAPS while answering. 2.It is always better not to give the final answer directly at first. Make sure you walk the asker to the answer instead of giving it away. Good luck!

OpenStudy (anonymous):

is redneck suspended?

OpenStudy (anonymous):

no, sir

OpenStudy (anonymous):

No, the question was towards the moderators, don't worry. It's all okay now : )

OpenStudy (anonymous):

I think this question is all over google now

OpenStudy (aravindg):

Trolling is not allowed. Let us see if he takes our advice or continues with random answers. Everyone should be given a chance :)